15. Ms. Joan Hanson is an employee of a Canadian controlled private corporation. During 2016, she receives options to purchase 500 shares of her employer’s common stock at a price of $22 per share. At this time, the estimated per share value of the stock is $20.50. During 2017, she exercises all of these options. At this time, the estimated market value of the stock is $31.50 per share. On December 1, 2017, she sells the stock for $38.75 per share. The net effect of the 2017 transactions on her Taxable Income would be: A. An increase of $1,812.50. B. An increase of $2,375.00. C. An increase of $4,187.50. D. An increase of $4,750.00. E. An increase of $6,562.50.
Answer #1
Employment Income = (Market Value - Exercise Price) X Number of Shares = (31.5 - 22) X 500 = $4,750
Capital Gain = (sale Price - Market price at exercise) X Number of shares = (38.75 - 31.5) X 500 = $3,625 50% = $1,812.5
Net Increase = 4,750 + 1,812.5 = $6,562.5
In 2017 SOR = (4,750 X 0.5) = $2,375
Net Increase = 6,562.5 - 2,375 = $4,187.5

Answer is option C